OpenStudy (anonymous):

[8.03] What are the factors of x2 + 12x + 20? (x + 4)(x + 5) (x + 2)(x + 10) (x − 4)(x − 5) (x − 2)(x − 10)

OpenStudy (anonymous):

u have to use this method first u should find the roots of the equation x1 and x2 r ur roots----->(x-x1)(x-x2)

OpenStudy (anonymous):

ax^2+bx+c=0 \[\Delta = b ^{2} -4ac\] \[x1=(-b+\sqrt{\Delta})/2a\] \[x2=(-b-\sqrt{\Delta})/2a\]